Andrew invests a total of 120,000 in two different accounts. the first account earns 5% while the second account ears 6% if the total interest earned after one year is 6700 what principal was invested in each account?
Answer by mananth(16946)

Investment I 5.00% per annum ---x
Investment II 6.00% per annum ---y

x + y= 120000 ------------------------1
5.00% x + 6.00% y= = $6,700.00
Multiply by 100
5 x + 6 y= = $670,000.00 --------2

Multiply (1) by -5
we get

-5 x -5 y= = -600000.00

Add this to (2)

0 x 1 y= = $70,000.00

divide by 1

y = $70,000.00 investment at 6.00%
Balance $50,000.00 investment at 5.00%
CHECK
$50,000.00 --------- 5.00% ------- $2,500.00
$70,000.00 ------- 6.00% ------- $4,200.00
Total -------------- $6,700.00